OpenStudy (anonymous):

Would someone mind finding the derivative of xy+x+y=5 please? My answer is different from what it should be

OpenStudy (anonymous):

xy + x + y - 5 = 0 (x + 1)y = 5 - x

OpenStudy (anonymous):

y+y'x+1+y' =0 y'(x+1) = -(y+1) y' = -(y+1)/(x+1)

OpenStudy (anonymous):

Is that what you're getting?

OpenStudy (anonymous):

@greenglasses No... :/ That was the right answer though. Thank you! I can see where i was wrong now :)

OpenStudy (anonymous):

Okay great :)